An annuity providing a rate of return of 4.8% compounded monthly was purchased for $45,000. The annuity pays $300 at the end of each month. What will be the principal portion of Payment 92? (Round your answer to 2 decimal places.)
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

SOLVED WITH BA II PLUS CALCULATOR

05:21 C w ENG 66 01-08-2020 35 Х E240 . А B с D E F G H - - K BA II PLUS SET C/Y=P/Y =12 218 219 220 221 222 223 224 4.8 1/Y